# Imaging Genetics Analysis of Alzheimer’s Disease

### v1

@@ Abstract
ordinal regression,
keywords:
doi:
0000
volume: 0issue: 0articletype: research-article
[…]
October
[type=corresp,id=t3]Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in the analysis or writing of this report. A complete listing of ADNI investigators can be found at: http://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
[type=corresp,id=c1]Corresponding author.
@@ 1. Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that impairs memory, thinking, and behavior, leading to severe cognitive decline over time. It is the most common cause of dementia, accounting for 60–80% of all dementia cases (Irwin et al., 2018). AD typically begins with mild memory loss and gradually progresses to more severe impairments, ultimately impacting a person’s ability to perform everyday tasks and maintain independence. The key pathological hallmarks of AD include the accumulation of amyloid-beta plaques and tau tangles in the brain, leading to neuronal damage and cognitive decline (DeTure and Dickson, 2019). Currently, there is no known cure, but early diagnosis and interventions can help manage symptoms and improve the quality of life for affected individuals.
@@
Cognitive impairment presents itself in several stages, beginning with Cognitively Normal (CN), where individuals do not display any significant cognitive issues. This is followed by Mild Cognitive Impairment (MCI), a transitional stage between normal aging and Alzheimer’s disease. MCI can be further categorized into early (EMCI) and late (LMCI) stages, based on the degree of impairment (Aisen et al., 2015). In its advanced form, cognitive decline progresses to AD, where patients experience significant memory loss, impaired reasoning, and behavioral disturbances. Neuropsychological exams are essential tools for assessing cognitive function and directly quantifying cognitive in diagnosed patients. These exams evaluate various domains such as memory, attention, executive function, and language, providing valuable insights for distinguishing normal aging, and differences between MCI versus AD (Nelson and O’Connor, 2008). Examples of these exams include the Alzheimer’s Disease Assessment Scores (ADAS), the Mini-Mental State Exam (MMSE), and more (Mohs and Cohen, 1988; Folstein and McHugh, 1975). As a commonly used assessment, the MMSE is scored on a 30-point scale, with higher scores indicating better cognitive function. Scores above 27 suggest normal cognition, scores between 24–27 indicate possible MCI, and scores below 24 typically indicate more severe impairments, such as in AD (Arevalo-Rodriguez, 2015).
@@
Dementia is associated with notable structural changes in the brain, including atrophy of the hippocampus and cortical regions responsible for memory and cognition (Wiseman et al., 2004). White matter integrity is often disrupted, and diffusion tensor imaging (DTI) studies have identified changes in Fractional Anisotropy (FA) (Kantarci et al., 2014; Mayo et al., 2019), a measure of white matter microstructure. Lower FA values reflect impaired neuronal connectivity, which is commonly observed in patients with AD and MCI (Tae et al., 2018). These structural abnormalities correspond to cognitive decline, making FA an important biomarker for studying disease progression. Genetic research plays a crucial role in understanding the development and progression of Alzheimer’s disease. Variants in genes such as apolipoprotein-E (APOE) allele (Troutwine et al., 2022) have been strongly associated with increased risk for AD, as shown in Figure 6 (Appendix A). Beyond APOE, genome-wide association studies (GWAS) have identified other genes involved in amyloid processing, neuroinflammation, and lipid metabolism that may contribute to AD pathology (Alzheimer Disease Genetics Consortium, ADGC). Exploring these genetic markers helps identify individuals at higher risk, informs personalized treatments, and advances the understanding of disease mechanisms.
@@ 2. ADNI-Data
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) is an ongoing longitudinal cohort study designed to develop clinical, imaging, and genetic markers for early detection and tracking of AD (Petersen, 2010). Eligible participants, aged 55-90 and in generally good health, were enrolled with either memory concerns or normal cognition. Each participant underwent comprehensive assessments, including cognitive testing, imaging, genetic evaluations, and both invasive and non-invasive medical procedures. Follow-up visits were conducted approximately every six months. The TADPOLE Challenge provided original data conducted from ADNI.
@@ 2.1 Cohort Description and Pre-Processing
Our study cohort consists of baseline visits from three phases: ADNI-1, ADNIGO, and ADNI-2 (Petersen, 2010; Weber et al., 2021; Beckett et al., 2015). All patients included in the analysis were de-identified and exhibited varying degrees of cognitive impairment. Although ADNI provides follow-up data, we focused exclusively on baseline data to better understand the population-level characteristics of cognitively impaired subjects. For data preprocessing, visualization, and statistical analysis, we utilized a combination of tools. Preprocessing tasks were conducted using Python 3, employing packages such as Numpy (Harris, 2020), Pandas (McKinney, 2010), and Matplotlib (Hunter, 2007) for data manipulation. For statistical analysis, we used R statistical software version 4.3.3. The primary goal of this study is to identify biomarkers associated with the acceleration of AD progression. After selecting baseline-only patients and removing incomplete or censored data, our final cohort consisted of 1,631 patients (819-ADNI1, 129-ADNIGO, and 683-ADNI2). Certain biological factors were excluded from analysis due to incomplete data or patient refusal. The selected predictors included are listed below in Table 1. A more comprehensive summary of biomarker statistics is provided in Appendix A Table 4.
@@
Our secondary objective is to investigate genetic associations with disease progression. We analyze gene expression data in relation to the biomarkers identified in our primary analysis and explored correlations between gene expression and brain structure changes by examining Fractional Anisotropic levels in diffusion tensor imaging (DTI) data. This analysis is insightful in identifying clusters of white matter associated with rapid neurodegeneration (Poulakis, 2021).
@@ 3.1.2 Ordinal Logistic Regression
The ordinal model is a statistical technique used to model the relationship between an ordinal dependent variable and independent variables. It utilizes logistic regression (like a link function) to measure the likelihood will fall into one of the categories. Our key assumption is that the relationship between predictors and log-odds of outcomes is proportional across all thresholds (proportional odds assumption.) (Yee, 2010)
@@ 3.2.1 Screening
Due to the vast amount of genetic information, we reduce the dimensionality. To achieve this, we applied the Sure Independence Screening method (Fan and Lv, 2008). SIS effectively selects a subset of the most important predictors by ranking them based on their marginal association with the outcome variable. By focusing on top-ranking features, SIS ensures that we do not lose important predictors, and we can then apply more refined variable selection techniques in a computationally efficient way.
We applied the Least Absolute Shrinkage and Selection Operator (LASSO) penalty-based regression methods. LASSO is particularly beneficial in our case because it automates model selection through its shrinkage effect, as well as preventing over-fitting (Tibshirani, 1996). It effectively reduces the coefficients of less significant genes to zero, concentrating on relevant, non-zero predictors.
@@ 3.2.2 Multivariate Response
The multi-response Gaussian family is particularly useful when there are several (correlated) responses, also known as the “multi-task learning” problem (Hastie et al., 2024). In this case, a covariate is either included in the model for all the responses or excluded for all the responses.
@@ 3.2.3 Multinomial Response
The multinomial model is particularly useful for classification problems where the response variable can take multiple categorical outcomes rather than just binary labels (Hastie et al., 2024). Unlike one-vs-all approaches, the multinomial model jointly models all class probabilities using the softmax function, ensuring that the probabilities sum to one. This allows the model to capture relationships between different categories rather than treating them independently.
@@ 4. Results
With three distinct analytical objectives, we begin with a full cohort of 1,631 patients, which we use for all low-dimensional analyses. From this cohort, we identify a subset of 468 patients with available genetic information. Among these, a further subset of 104 patients also has Fractional Anisotropy (FA) imaging data available. This sub-subset is used for imaging-informed kernel-based analyses. Across all levels of analysis, a p-value threshold of <0.05<0.05 is considered statistically significant. We apply the following R packages for our analysis: glmnet (Friedman et al., 2010; Simon et al., 2011; Tay et al., 2023), SIS (Saldana and Feng, 2018), corrplot (Wei and Simko, 2021), ggplot2 (Wickham, 2016), MASS (Venables and Ripley, 2002), car (Fox and Weisberg, 2019), and gtsummary (Sjoberg et al., 2021).
@@ 4.1.1 Regressing Disease State on Cognitive Scores
The results highlight statistically significant associations between several cognitive exam scores and disease progression, as visualized in Figure 1. Among the most influential predictors, the Mini-Mental State Examination (MMSE) and the Clinical Dementia Rating–Sum of Boxes (CDRSB) display the smallest and largest odds ratios, respectively. This finding is consistent with clinical expectations, as lower MMSE scores and higher CDRSB scores reflect worsening cognitive function and more advanced stages of Alzheimer’s disease (Andrews et al., 2019; Lanctôt et al., 2024). In contrast, APOE4 carrier status shows a positive and statistically significant association with disease progression (Qian et al., 2021). Variables such as education level, baseline age, and RAVLT (Rey Auditory Verbal Learning Test) scores do not reach statistical significance (p-values > 0.05) in the multivariate model. Notably, MMSE and CDRSB scores demonstrate extreme magnitudes, indicating that each per-unit increase in MMSE provides a protective effect, while increments in CDRSB signal worsening cognitive impairment. Further analysis is required to investigate these biomarkers.
@@ 4.2.1 Regressing CS on Gene Expression
Recognizing the importance of MMSE and CDRSB as key markers of cognitive decline, we examine their associations from a genetic perspective. To address the high dimensionality of the genetic data, we apply a marginal screening technique using the Sure Independence Screening method (Mishra and Li, 2020), which ranks all predictors and retains only the top variables for further analysis. We use our genetic subsample (n = 468), we perform SIS with the default setting of 10 partitions (nfolds = 10). We then fit a linear model with a Gaussian family, regressing the selected cognitive biomarkers (MMSE and CDRSB) on the screened genetic features. Only genes with non-zero coefficients in this final model are considered as potential predictors of cognitive decline.
@@
Figure 2 illustrates the top 75 genes identified based on the estimated regression coefficients, with the size of each gene name proportional to the magnitude of the coefficients in relation to cognitive function. For MMSE and CDRSB as responses to capture cognitive decline across disease stages, four key genes—NAB2, 5-MAR, TGFBR1 (Bosco et al., 2013), and CLIC1—emerge as statistically significant in both models, with CLIC1 (Carlini, 2020) showing the greatest effect.
@@ 4.2.2 Regressing Disease State on Gene Expression
We repeated the analysis to identify key genes across disease progression stages. Since SIS does not support multinomial models, we apply SIS, treating each pairwise comparison between disease stages (CN, EMCI, LMCI, AD) as separate binary classifications (Hasan Mahmud et al., 2023; Sener et al., 2024). After SIS screening, we applied cross-validation using a multinomial model to capture genetic effects varying by disease stage. This approach allowed detection of genes significant in some transitions but not others. Finally, we compiled union and intersection sets of genes with non-zero coefficients across comparisons.
@@
Figure 3 highlights the associations between intersecting genes and cognitive states, with color gradients indicating effect magnitude. Notably, CCNI (Hirabayashi et al., 2025) shows a strong positive association with Alzheimer’s Disease, while PRNP has moderate associations in early to late cognitive decline stages. In contrast, genes like YIPF6 and RAB27A show negligible associations. These findings suggest that a gene which shows high effect at earlier or later stages of impairment may display a null effect at different stages. As such, finding intersections of the predictors reinforces our notion that a significant effect occurs in underlying mechanisms (Zakeri and et al., 2020). In relation to the cognitive exam analyses from the previous step, we note that the following genes—TGFBR1, XRCC5, and CCNI—are correlated with both disease progression and individual cognitive scorings. It is important to note that genetic effects may not reflect a constant effect as disease state progresses.
@@ 4.2.3 Imaging Genetic Analysis
For our image analysis, we categorize the brain sections into three groups: the Corpus Callosum (CC), Right Brain, and Left Brain. The CC connects the two cerebral hemispheres and is crucial for higher cognitive functions. It is often one of the first regions to show signs of neurodegeneration. The left hemisphere handles language processing, analytical thinking, and logic, while the right hemisphere is more involved in spatial awareness and emotional interpretation. Damage to these structures can lead to disruptions in communication between the hemispheres, progressing neurodegenerative mechanisms (Dorion et al., 2022; Zhang et al., 2024).
@@
Figure 5 indicates the top 50 genes selected from the multi-Gaussian fitting process, with the size of each gene name reflecting its significance in relation to white matter integrity. Using the FA measurements to capture genetic effect, many predictors emerge as statistically significant in all three brain regions. Some of the listed genes include CTSB (Hook et al., 2023), TMEM184A, POLR2D (Chum et al., 2024), INVS, ANXA4 (White et al., 2024), and others.
Our results suggest that the top genes in the two hemispheres share more commonalities compared to those in the corpus callosum. For instance, ENTPD6 is identified as significant in measuring white matter in the hemispheres but shows no significance in the CC. This finding may indicate that the genetic mechanisms governing white matter in the hemispheres are more uniform or shared, reflecting their complementary roles in processing information and supporting localized brain functions (Zhao et al., 2021). In contrast, the gene TMEM184A demonstrates significance across all three regions, though with weaker effect sizes. Genes like TMEM184A, which are consistently significant across multiple brain regions, can serve as stronger indicators of white matter degradation and may hold greater potential as predictive markers.

### v2

@@ 1. Introduction
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that impairs memory, thinking, and behavior, leading to severe cognitive decline over time. It is the most common cause of dementia, accounting for 60–80% of all dementia cases [1]. AD typically begins with mild memory loss and gradually progresses to more severe impairments, ultimately impacting a person’s ability to perform everyday tasks and maintain independence. The key pathological hallmarks of AD include the accumulation of amyloid-beta plaques and tau tangles in the brain, leading to neuronal damage and cognitive decline [2]. Currently, there is no known cure, but early diagnosis and interventions can help manage symptoms and improve the quality of life for affected individuals.
@@
Cognitive impairment presents itself in several stages, beginning with Cognitively Normal (CN), where individuals do not display any significant cognitive issues. This is followed by Mild Cognitive Impairment (MCI), a transitional stage between normal aging and Alzheimer’s disease. MCI can be further categorized into early (EMCI) and late (LMCI) stages, based on the degree of impairment [3]. In its advanced form, cognitive decline progresses to AD, where patients experience significant memory loss, impaired reasoning, and behavioral disturbances. Neuropsychological exams are essential tools for assessing cognitive function and directly quantifying cognitive in diagnosed patients. These exams evaluate various domains such as memory, attention, executive function, and language, providing valuable insights for distinguishing normal aging, and differences between MCI versus AD [4]. Examples of these exams include the Alzheimer’s Disease Assessment Scores (ADAS), the Mini-Mental State Exam (MMSE), and more [5, 6]. As a commonly used assessment, the MMSE is scored on a 30-point scale, with higher scores indicating better cognitive function. Scores above 27 suggest normal cognition, scores between 24–27 indicate possible MCI, and scores below 24 typically indicate more severe impairments, such as in AD [7].
@@
Dementia is associated with notable structural changes in the brain, including atrophy of the hippocampus and cortical regions responsible for memory and cognition [8]. White matter integrity is often disrupted, and diffusion tensor imaging (DTI) studies have identified changes in Fractional Anisotropy (FA) [9, 10], a measure of white matter microstructure. Lower FA values reflect impaired neuronal connectivity, which is commonly observed in patients with AD and MCI [11]. These structural abnormalities correspond to cognitive decline, making FA an important biomarker for studying disease progression. Genetic research plays a crucial role in understanding the development and progression of Alzheimer’s disease. Variants in genes such as apolipoprotein-E (APOE) allele [12] have been strongly associated with increased risk for AD, as shown in Figure 6 (Appendix A). Beyond APOE, genome-wide association studies (GWAS) have identified other genes involved in amyloid processing, neuroinflammation, and lipid metabolism that may contribute to AD pathology [13]. Exploring these genetic markers helps identify individuals at higher risk, informs personalized treatments, and advances the understanding of disease mechanisms.
@@ 2. ADNI-Data
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) is an ongoing longitudinal cohort study designed to develop clinical, imaging, and genetic markers for early detection and tracking of AD [14]. Eligible participants, aged 55-90 and in generally good health, were enrolled with either memory concerns or normal cognition. Each participant underwent comprehensive assessments, including cognitive testing, imaging, genetic evaluations, and both invasive and non-invasive medical procedures. Follow-up visits were conducted approximately every six months. The TADPOLE Challenge provided original data conducted from ADNI.
@@ 2.1 Cohort Description and Pre-Processing
Our study cohort consists of baseline visits from three phases: ADNI-1, ADNIGO, and ADNI-2 [14, 15, 16]. All patients included in the analysis were de-identified and exhibited varying degrees of cognitive impairment. Although ADNI provides follow-up data, we focused exclusively on baseline data to better understand the population-level characteristics of cognitively impaired subjects. For data preprocessing, visualization, and statistical analysis, we utilized a combination of tools. Preprocessing tasks were conducted using Python 3, employing packages such as Numpy [17], Pandas [18], and Matplotlib [19] for data manipulation. For statistical analysis, we used R statistical software version 4.3.3. The primary goal of this study is to identify biomarkers associated with the acceleration of AD progression. After selecting baseline-only patients and removing incomplete or censored data, our final cohort consisted of 1,631 patients (819-ADNI1, 129-ADNIGO, and 683-ADNI2). Certain biological factors were excluded from analysis due to incomplete data or patient refusal. The selected predictors included are listed below in Table 1. A more comprehensive summary of biomarker statistics is provided in Appendix A Table 4.
@@
Our secondary objective is to investigate genetic associations with disease progression. We analyze gene expression data in relation to the biomarkers identified in our primary analysis and explored correlations between gene expression and brain structure changes by examining Fractional Anisotropic levels in diffusion tensor imaging (DTI) data. This analysis is insightful in identifying clusters of white matter associated with rapid neurodegeneration [20].
@@ 3.1.2 Ordinal Logistic Regression
The ordinal model is a statistical technique used to model the relationship between an ordinal dependent variable and independent variables. It utilizes logistic regression (like a link function) to measure the likelihood will fall into one of the categories. Our key assumption is that the relationship between predictors and log-odds of outcomes is proportional across all thresholds (proportional odds assumption.) [21]
@@ 3.2.1 Screening
Due to the vast amount of genetic information, we reduce the dimensionality. To achieve this, we applied the Sure Independence Screening method [22]. SIS effectively selects a subset of the most important predictors by ranking them based on their marginal association with the outcome variable. By focusing on top-ranking features, SIS ensures that we do not lose important predictors, and we can then apply more refined variable selection techniques in a computationally efficient way.
We applied the Least Absolute Shrinkage and Selection Operator (LASSO) penalty-based regression methods. LASSO is particularly beneficial in our case because it automates model selection through its shrinkage effect, as well as preventing over-fitting [23]. It effectively reduces the coefficients of less significant genes to zero, concentrating on relevant, non-zero predictors.
@@ 3.2.2 Multivariate Response
The multi-response Gaussian family is particularly useful when there are several (correlated) responses, also known as the “multi-task learning” problem [24]. In this case, a covariate is either included in the model for all the responses or excluded for all the responses.
@@ 3.2.3 Multinomial Response
The multinomial model is particularly useful for classification problems where the response variable can take multiple categorical outcomes rather than just binary labels [24]. Unlike one-vs-all approaches, the multinomial model jointly models all class probabilities using the softmax function, ensuring that the probabilities sum to one. This allows the model to capture relationships between different categories rather than treating them independently.
@@ 4. Results
With three distinct analytical objectives, we begin with a full cohort of 1,631 patients, which we use for all low-dimensional analyses. From this cohort, we identify a subset of 468 patients with available genetic information. Among these, a further subset of 104 patients also has Fractional Anisotropy (FA) imaging data available. This sub-subset is used for imaging-informed kernel-based analyses. Across all levels of analysis, a p-value threshold of <0.05<0.05 is considered statistically significant. We apply the following R packages for our analysis: glmnet [25, 26, 27], SIS [28], corrplot [29], ggplot2 [30], MASS [31], car [32], and gtsummary [33].
@@ 4.1.1 Regressing Disease State on Cognitive Scores
The results highlight statistically significant associations between several cognitive exam scores and disease progression, as visualized in Figure 1. Among the most influential predictors, the Mini-Mental State Examination (MMSE) and the Clinical Dementia Rating–Sum of Boxes (CDRSB) display the smallest and largest odds ratios, respectively. This finding is consistent with clinical expectations, as lower MMSE scores and higher CDRSB scores reflect worsening cognitive function and more advanced stages of Alzheimer’s disease [34, 35]. In contrast, APOE4 carrier status shows a positive and statistically significant association with disease progression [36]. Variables such as education level, baseline age, and RAVLT (Rey Auditory Verbal Learning Test) scores do not reach statistical significance (p-values >> 0.05) in the multivariate model. Notably, MMSE and CDRSB scores demonstrate extreme magnitudes, indicating that each per-unit increase in MMSE provides a protective effect, while increments in CDRSB signal worsening cognitive impairment. Further analysis is required to investigate these biomarkers.
@@ 4.2.1 Regressing CS on Gene Expression
Recognizing the importance of MMSE and CDRSB as key markers of cognitive decline, we examine their associations from a genetic perspective. To address the high dimensionality of the genetic data, we apply a marginal screening technique using the Sure Independence Screening method [37], which ranks all predictors and retains only the top variables for further analysis. We use our genetic subsample (n = 468), we perform SIS with the default setting of 10 partitions (nfolds = 10). We then fit a linear model with a Gaussian family, regressing the selected cognitive biomarkers (MMSE and CDRSB) on the screened genetic features. Only genes with non-zero coefficients in this final model are considered as potential predictors of cognitive decline.
@@
Figure 2 illustrates the top 75 genes identified based on the estimated regression coefficients, with the size of each gene name proportional to the magnitude of the coefficients in relation to cognitive function. For MMSE and CDRSB as responses to capture cognitive decline across disease stages, four key genes—NAB2, 5-MAR, TGFBR1 [38], and CLIC1—emerge as statistically significant in both models, with CLIC1 [39] showing the greatest effect.
@@ 4.2.2 Regressing Disease State on Gene Expression
We repeated the analysis to identify key genes across disease progression stages. Since SIS does not support multinomial models, we apply SIS, treating each pairwise comparison between disease stages (CN, EMCI, LMCI, AD) as separate binary classifications [40, 41]. After SIS screening, we applied cross-validation using a multinomial model to capture genetic effects varying by disease stage. This approach allowed detection of genes significant in some transitions but not others. Finally, we compiled union and intersection sets of genes with non-zero coefficients across comparisons.
@@
Figure 3 highlights the associations between intersecting genes and cognitive states, with color gradients indicating effect magnitude. Notably, CCNI [42] shows a strong positive association with Alzheimer’s Disease, while PRNP has moderate associations in early to late cognitive decline stages. In contrast, genes like YIPF6 and RAB27A show negligible associations. These findings suggest that a gene which shows high effect at earlier or later stages of impairment may display a null effect at different stages. As such, finding intersections of the predictors reinforces our notion that a significant effect occurs in underlying mechanisms [43]. In relation to the cognitive exam analyses from the previous step, we note that the following genes—TGFBR1, XRCC5, and CCNI—are correlated with both disease progression and individual cognitive scorings. It is important to note that genetic effects may not reflect a constant effect as disease state progresses.
@@ 4.2.3 Imaging Genetic Analysis
For our image analysis, we categorize the brain sections into three groups: the Corpus Callosum (CC), Right Brain, and Left Brain. The CC connects the two cerebral hemispheres and is crucial for higher cognitive functions. It is often one of the first regions to show signs of neurodegeneration. The left hemisphere handles language processing, analytical thinking, and logic, while the right hemisphere is more involved in spatial awareness and emotional interpretation. Damage to these structures can lead to disruptions in communication between the hemispheres, progressing neurodegenerative mechanisms [44, 45].
@@
Figure 5 indicates the top 50 genes selected from the multi-Gaussian fitting process, with the size of each gene name reflecting its significance in relation to white matter integrity. Using the FA measurements to capture genetic effect, many predictors emerge as statistically significant in all three brain regions. Some of the listed genes include CTSB [46], TMEM184A, POLR2D [47], INVS, ANXA4 [48], and others.
Our results suggest that the top genes in the two hemispheres share more commonalities compared to those in the corpus callosum. For instance, ENTPD6 is identified as significant in measuring white matter in the hemispheres but shows no significance in the CC. This finding may indicate that the genetic mechanisms governing white matter in the hemispheres are more uniform or shared, reflecting their complementary roles in processing information and supporting localized brain functions [49]. In contrast, the gene TMEM184A demonstrates significance across all three regions, though with weaker effect sizes. Genes like TMEM184A, which are consistently significant across multiple brain regions, can serve as stronger indicators of white matter degradation and may hold greater potential as predictive markers.